# Forgetting-MarI: LLM Unlearning via Marginal Information Regularization

### v1

@@ 1 Introduction
We argue that the former objective is the natural objective when people talk about unlearning. Moreover, that LLM unlearning naturally targets marginal unlearning, and that existing LLM unlearning methods indirectly target it by balancing signals from data to unlearn and data to retain. (see Appendix A.1). Indeed, the goal of unlearning is not to eradicate knowledge contained in the unlearn data, but rather to surgically remove only its marginal effect, the information not already supported by the data we are authorized to use. In this copyright scenario, the marginal effect unlearning satisfies legal requirements with minimal utility loss, whereas the full removal would unnecessarily discard information that is lawfully present in the model. This distinction motivates our proposed method, Forgetting-MarI, a direct marginal information***The term Marginal Information is formalized in Definition 1.1. removal of the unlearned data.
Figure 1 further demonstrates the difference between full-information unlearning and marginal-information unlearning (detailed experimental setup in Section 4.3). We created three models trained on ground truth prompts: one before unlearning, one after marginal information unlearning, and one after full information unlearning. Models are given the first half of a sentence (prompt) and are asked to complete it. Before unlearning, the model completes the sentences in a way that is similar to the ground truth. With marginal unlearning, the model produces different but coherent completions. With full unlearning, the model struggles to coherently complete the sentences.
@@ 1.2 Our Contributions
Definition 1.1 (Marginal Information (MarI)).
@@
  • (A4) Exemplary experimental performance: Experiments show that our proposed method outperforms state-of-the-art unlearning methods in a wide range of unlearning tasks.
@@ 2.1 Quantifying and Unlearning Marginal Effects
Proposition 2.1 (Detection accuracy upper bounded by mutual information).
@@
Definition 2.1 (MI-based marginal information loss).
@@
where $\ell_{KL}(\theta,r):=D_{\mathrm{KL}}\!\big(p^{\,r}(\theta)\,\|\,p^{\,r}(\theta_{0})\big)$ is the KL divergence between the updated model (parameter $\theta$) and the frozen original model (parameter $\theta_{0}$) on $r$, and $\ell_{\operatorname{MarI}}$ is as above, motivated by Prop. 2.1. Algorithm 3 describes an efficient LLM implementation.
Remark 2.1 (Alternative quantification).
@@ 2.2 Marginal Information & Perplexity-Based Detectors
Theorem 2.1 (MarI controls the self-perplexity gap).
@@
Finally, we show that MarI directly controls the score gap even for a neighborhood of $u$ rather than only $u$ itself. Proof can be found in Appendix B.4.
Theorem 2.2 (MarI controls neighborhood-perplexity gap).
Draw $U:=\{U_{t}\}_{t=1}^{T}$ with $U_{t}\sim p_{t}^{\,u}$ independently across $t\in[T]$ and suppose $\max_{t,x}\,\big\{\frac{p_{t}^{\,u}(x)}{p_{t}^{\,r}(x)}\!\vee\!\frac{p_{t}^{\,r}(x)}{p_{t}^{\,u}(x)}\big\}\leq M$. Let $C:=\max_{t,x:\,p_{t}^{\,u}(x)>0}\bigl[\log\frac{p_{t}^{\,r}(x)}{p_{t}^{\,u}(x)}\bigr]^{2}<\infty$. Then, for any ${\varepsilon}>0$, with probability at least $1-2\exp\bigl(-T{\varepsilon}^{2}/(2C)\bigr)$,
@@ 3 Algorithm Design
Algorithm 3 presents pseudo-code for Forgetting-MarI. In theory, the token-wise MarI (Equation 1) provides strong guarantees when sentences in $r$ and $u$ are homogeneous in length and token-wise context. In practice, however, token-wise MarI Loss can be noisy under heterogeneous batches. To address this, we also provide a pooled (“flattened”) estimator that first averages across token positions (and batch) to form $\bar{p}^{\,s}=\frac{1}{T}\sum_{t}p_{t}^{\,s},\,s\in\{r,u,d\}$, then computes the pooled MarI Loss $I(\bar{X}_{\operatorname{MarI}};Z)=\operatorname{\mathrm{JSD}}\,\bigl(\bar{p}^{\,d},\bar{p}^{\,r}\bigr)$. By the data-processing inequality, $I(\bar{X}_{\operatorname{MarI}};Z)$ is a variational lower bound to the token-wise MarI, $I(X_{\operatorname{MarI}};Z)$. Such a pooled version aims to stabilize marginal information quantification by filtering the position-heterogeneous noise and emphasizing the dominant distribution shift.
Algorithm 1 Forgetting-MarI.
In practice, one should use hetero=False when $r$ and $u$ are homogeneous (similar lengths and aligned token-wise contexts), to leverage the more accurate token-wise signal. Use hetero=True for large dataset with random batching and unaligned sequences, to obtain a more stable pooled marginal information signal.
A full theoretical/empirical comparison (and mixtures of both signals) is deferred to future work. In our experiments, $I(X_{\operatorname{MarI}};Z)$ and pooled $I(\bar{X}_{\operatorname{MarI}};Z)$ achieve comparable unlearning under different hyperparameters. Therefore, Section 4 reports results under one label. We include a head-to-head comparison between the two approaches of marginal information in Appendix C.
@@ 4 Experiments
Table 3: The table summarizes the utility and unlearning types of methods included in the comparison experiments.
We evaluated Forgetting-MarI on GPT and Llama family models on two real-world datasets: Harry Potter and the Prisoner of Azkaban [rowling2013harry] and Careless People: A Cautionary Tale of Power, Greed, and Lost Idealism [wynn2025careless], chosen for contrasting genre and pretraining prevalence (Harry Potter is abundant in pretraining; Careless People is likely scarce).GPT experiments were run on a node with four NVIDIA A100 GPUs (80 GB HBM2 each), while Llama experiments were run using four NVIDIA T4 GPUs (16 GB GDDR6 each) As a full-parameter unlearning approach, we benchmark against full-parameter fine-tuning baselines: gradient ascent (GA) [yao2024llmunlearning], KL-gradient ascent (KL-GA) [hong2024dissect], gradient difference (GD) [liu2022continual], and direct preference optimization (DPO) [rafailov2023dpo]. Table 3 summarizes these baselines alongside other unlearning approaches. Detailed discussion of the latter is deferred to Appendix A.1.
Our numerical results demonstrate four key advantages: (1) Balanced unlearning: after hyperparameter tuning for all methods, Forgetting-MarI best matches the retain–unlearn trade-off, preserving performance on $\mathcal{D}_{r}$ while removing $\mathcal{D}_{u}$; (2) Stable continual unlearning: in terms of both regularization parameter selection and model performance during training, Forgetting-MarI is the most stable; (3) General capacity: models unlearned with Forgetting-MarI retain general model capabilities; (4) Theoretical guarantee: perplexity/confidence-based detectors can no longer detect unlearning data following unlearning via Forgetting-MarI, as guaranteed by Theorems 2.2 and 2.1.
4.1 Experimental Setup
To simulate real-world scenarios where forget and retain content is semantically intertwined, we construct the unlearn set $\mathcal{D}_{u}$ and retain set $\mathcal{D}_{r}$ by selecting alternating sentences from each dataset. We use a separate validation set to monitor utility. For Careless People, we use Reddit stories for this validation set [writingprompts]. Experiments share the following protocol: (i) fine-tune on $\mathcal{D}_{u}\!\cup\!\mathcal{D}_{r}$ to obtain a baseline; (ii) apply each unlearning method to remove the influence of $\mathcal{D}_{u}$; (iii) train a gold-standard unlearn baseline on $\mathcal{D}_{r}$ only, simulating a model that never saw $\mathcal{D}_{u}$. An optimally unlearned model should match the performance of this gold-standard. Model performance on the unlearn, retain, and validation sets were assessed using next-token prediction accuracy. General model performance was assessed using benchmarks from Eleuther’s LM Evaluation Harness [eval-harness].
Figure 2: Next-token prediction accuracies for different unlearning methods compared to the baselines.
4.2 Forgetting Harry Potter
Setting We fine-tune a pretrained GPT-2 Large [radford2019language] model on excerpts of Harry Potter and the Prisoner of Azkaban [rowling2013harry] following a similar approach of Eldan et al. [eldan2023approximate]. We curated equally-sized (50/50) retain and unlearn datasets with high contextual overlap to maximize their correlation.
Utility Preservation Figure 2 reports the best accuracies achieved for each unlearning method after tuning their regularization parameters (e.g., with the best $\lambda$ marked below the methods). The reported accuracies correspond to the training epoch when validation accuracy fell by 3% or more from the starting accuracy. Forgetting-MarI shows the best balance of unlearning and utility preservation, as it best matches the retain-unlearn trade-off of the gold-standard unlearn baseline.
Stability of continual unlearning Figure 3 highlights two notions of robustness:
Figure 3: Training curves for each method with varying choices of the regularization parameter $\lambda$. Forgetting-MarI exhibits smooth monotone behavior, while the other methods show oscillation or utility collapse.
(1) Training robustness over epochs. Forgetting-MarI descends steadily to its optimum. GA and GD overshoot and bounce, KL–GA diverges after 5–6 epochs, DPO plateaus prematurely.
(2) Robustness against regularization tuning. Forgetting-MarI shows a monotone and smooth utility-unlearning trade-off when adjusting the regularization parameter. In comparison, GD and KL-GA display unstable oscillations with different choices of $\lambda$.
Both training and regularization robustness are necessary for a practical use of unlearning techniques. Practitioners do not have access to ground truth baselines and have limited time to select/determine the best parameter or training epoch to stop at, so stability is essential. Forgetting-MarI is the most stable technique during unlearning, making it the safest choice in practice. We refer the reader to Appendix D Figure 10 for a smaller learning rate training curve, which further demonstrates the stability of continual unlearning.
Table 4: Performance on the WikiText benchmark.
Table 5: Performance on the MMLU hierarchical evaluation.
Preserved general model capacity To assess the general model capacity after unlearning, we compare the unlearned model with the baseline, the unlearn baseline, and GPT2-LG (the model before finetuning on the retain set) on the following benchmarks: WikiText and MMLU [eval-harness]. We repeated the experiment with a realistic 10/90 unlearn/retain split to simulate practical scenarios where only a small portion of the data requires removal. The training curves for this setup are in Figure 9. Forgetting-MarI exhibits small declines on a few capability metrics (Tables 4 and 5) compared with the baseline or the unlearn baseline, while matching them on most tested metrics. Overall, the method largely preserves general model capacity on standard evaluations. Additional results on PIQA, ARC, and HellaSwag are reported in Table 7 (Appendix D).
4.3 Forgetting Careless People
Setting We fine-tune Meta’s Llama-3.2-1B [huggingfaceMetallamaLlama321BHugging] on Careless People: A Cautionary Tale of Power, Greed, and Lost Idealism [wynn2025careless], following the procedure outlined in [maini2024tofutaskfictitiousunlearning]. Beyond the setup in Sec. 4.1, we also test an uncorrelated scenario, where we used 2025 Reddit stories as $\mathcal{D}_{u}$ and the other half of Careless People as validation. Both datasets were published after the Llama model’s release, ensuring $\mathcal{D}_{u}$ is not in the model’s pre-training.
Utility preservation Figure 4 shows the best results for each unlearning method on both experiments. Consistent with results in Sec. 4.2, Forgetting-MarI most closely matches the unlearn baseline, achieving comparable unlearn accuracy while maintaining a similar retain and validation accuracy in both settings:
(i) Correlated experiment: GD and DPO both overtrain on $\mathcal{D}_{r}$, while KL-GA struggles to remove the information in $\mathcal{D}_{u}$ while maintaining performance on $\mathcal{D}_{r}$. Moreover, all other methods show a gradual loss of performance on the test set. (ii) Uncorrelated experiment: With less informational overlap between $\mathcal{D}_{u}$ and $\mathcal{D}_{r}$, other unlearning methods more easily remove $\mathcal{D}_{u}$. However, GD and DPO still overtrain on $\mathcal{D}_{r}$ while either over- or undershooting the desired unlearning. KL-GA quickly overshoots the desired amount of unlearning.
Figure 4: Next-token prediction accuracy for correlated (top) and uncorrelated (bottom) test sets.
Stability of continual unlearning Figure 5 shows the unlearning performance of each method over the course of unlearning, where the curves for each method correspond to the experiment with the best performing regularization parameter for each method.
Figure 5: Next-token prediction accuracy during training for both experiments. Epoch 0 represents the model before unlearning. Horizontal lines represent the “gold standard” unlearn baseline model, the model trained only on $\mathcal{D}_{r}$. Left/right columns for each method show correlated/uncorrelated test results respectively.
Table 6: Evaluation results from the ARC-Easy and PIQA benchmarks for both the correlated and uncorrelated experiments.
We note that there is the possibility that, in theory, one could find a perfect balance between gradient ascent and utility regularization, leading to a stable balance between unlearning and utility preservation, using one of the other methods. However, such a balance seems practically unattainable due to the unlearning instability over time and the lack of monotonicity in the choice of $\lambda$ for methods based on gradient ascent.
Preserved general model capacity Table 6 shows the performance of Llama-3.2-1B unlearned with Forgetting-MarI on the ARC-Easy and PIQA [eval-harness] benchmarks. Similar to Sec. 4.2, the model maintains relatively little performance drop compared to the baseline and the unlearn baseline, with the largest gap being observed on the ARC-Easy benchmark for the uncorrelated experiment.
4.4 Empirical verification of theoretical guarantees

### v2

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
We argue that the former objective is the natural objective when people talk about unlearning and LLM unlearning naturally targets marginal unlearning. Indeed, the goal of unlearning is not to eradicate knowledge contained in the unlearn data, but rather to surgically remove only its marginal effect, the information not already supported by the data we are authorized to use. In this copyright scenario, the marginal effect unlearning satisfies legal requirements with minimal utility loss, whereas the full removal would unnecessarily discard information that is lawfully present in the model.
This distinction motivates our proposed method, Forgetting-MarI, a direct marginal information***The term Marginal Information is formalized in Definition 1.1. removal of the unlearned data. More specifically, marginal information unlearning optimizes an objective that directly measures and suppresses only the additional information contributed by the unlearn set beyond what is already supported by the retain set, where the utility term only aims to stabilize retain performance or help the model learn new datasets if needed. There is no intrinsic conflict between the unlearn and utility objectives. In contrast, existing LLM unlearning methods are full-information in principle: their ascent or preference loss term targets the entire signal of $\mathcal{D}_{u}$ (e.g., maximizing CE on $\mathcal{D}_{u}$) and they attempt to indirectly spare shared/legitimate knowledge by counterbalancing this with a retain loss (CE or KL), preference shaping, parameter subtraction, or orthogonality (see details Appendix A.1). In other words, there is an intrinsic conflict between the utility and unlearning objective, which is necessary for the counterbalance to work, but often leads to unstable unlearning and requires extreme effort in parameter-tuning.
Figure 1 further demonstrates the difference between full-information unlearning and marginal-information unlearning (detailed experimental setup in Section 4). We created three models trained on ground truth prompts: one before unlearning, one after marginal information unlearning, and one after full information unlearning. Models are given the first half of a sentence (prompt) and are asked to complete it. Before unlearning, the model completes the sentences in a way that is similar to the ground truth. With marginal unlearning, the model produces different but coherent completions. With full unlearning, the model struggles to coherently complete the sentences.
@@ 1.2 Our Contributions
Definition 1.1 (Marginal Information (MarI))
@@
  • (A4) Exemplary experimental performance: Experiments show that our proposed method outperforms state-of-the-art unlearning methods in unlearning tasks using real-world text data on mid-scale LLMs.
@@ 2.1 Quantifying and Unlearning Marginal Effects
Proposition 2.1 (Detection accuracy upper bounded by mutual information)
@@
Definition 2.1 (MI-based marginal information loss)
@@
where $\ell_{KL}(\theta,r):=D_{\mathrm{KL}}\!\big(p^{\,r}(\theta)\,\|\,p^{\,r}(\theta_{0})\big)$ is the KL divergence between the updated model (parameter $\theta$) and the frozen original model (parameter $\theta_{0}$) on $r$, and $\ell_{\operatorname{MarI}}$ is as above, motivated by Prop. 2.1. Algorithm D.2 describes an efficient LLM implementation.
Remark 2.1 (Alternative quantification)
@@ 2.2 Marginal Information & Perplexity-Based Detectors
Theorem 2.1 (MarI controls the self-perplexity gap)
@@
Finally, we show that MarI directly controls the score gap even for a neighborhood of $u$ rather than only $u$ itself. The formal result and its proof can be found in Appendix B.4.
@@ 3 Algorithm Design
Token-wise MarI (Equation 1) provides strong guarantees when sentences in $r$ and $u$ are homogeneous in length and token-wise context. In practice, however, token-wise MarI Loss can be noisy under heterogeneous batches. To address this, we also provide a pooled (“flattened”) estimator that first averages across token positions (and batch) to form $\bar{p}^{\,s}=\frac{1}{T}\sum_{t}p_{t}^{\,s},\,s\in\{r,u,d\}$, then computes the pooled MarI Loss $I(\bar{X}_{\operatorname{MarI}};Z)=\operatorname{\mathrm{JSD}}\,\bigl(\bar{p}^{\,d},\bar{p}^{\,r}\bigr)$. Such a pooled version aims to stabilize the marginal information quantification by filtering the position-heterogeneous noise and emphasizing the dominant distribution shift.
By the data-processing inequality, the pooled estimator $I(\bar{X}_{\operatorname{MarI}};Z)$ is a variational lower bound to the token-wise MarI. The gap between the pooled and token-wise MarI losses is controlled by the $\ell^{2}$-deviation of the token-sequence densities (details in Theorem C.1[Pooling error bound], Appendix C). Furthermore, pooled MarI offers a specific word-wise guarantee:
Theorem 3.1 (Word-level provable unlearning via pooled MarI)
Fix a token $w\in V$ and assume $\bar{p}^{\,u}(w)\wedge\bar{p}^{\,r}(w)=:\bar{\gamma}_{w}\in(0,1]$. Then
$\Big|\,\log\bar{p}^{\,u}(w)-\log\bar{p}^{\,r}(w)\,\Big|\;\leq\;\frac{2\sqrt{2}}{\bar{\gamma}_{w}\,(1-\alpha)}\;\sqrt{\,I(\bar{X}_{\operatorname{MarI}};Z)\,}.$
Proof in Appendix C. Although weaker than the sentence-wise guarantee (Theorem 2.1), this guarantee is sufficient for provable unlearning of specific tokens and can be useful for the provable removal of word-level (not related to sentence structure) information.
Figure 2: Pseudo-code for Forgetting-MarI.
Figure 2 presents pseudo-code for Forgetting-MarI with pooled MarI. A full, detailed algorithm is provided as Figure D.2 in Appendix C. We recommend the following protocol:
  • Estimator Selection: Use token-wise MarI for homogeneous data (aligned contexts) to leverage precise signals. Use pooled MarI for large corpora with random batching to ensure stability.
  • Hyperparameter Tuning: Fix a tolerable perplexity gap for the application, derive the required MarI threshold, and tune $\gamma$ until the regularization magnitude falls below this threshold. This invokes the guarantees of Theorem 2.1 or 3.1.
As both estimators achieve comparable empirical results, Section 4 reports findings under a unified label (comparative ablation in Appendix C).
@@ 4 Experiments
Our experiments are designed to address three questions:
  • 1
    Utility-Forgetting Trade-off: Does Forgetting-MarI balances performance preservation on $\mathcal{D}_{r}$ while removing $\mathcal{D}_{u}$? By achieving similar unlearn performance $\mathcal{D}_{u}$ and utility preservation on $\mathcal{D}_{r}$ to the unlearn baseline (i.e., the retrain on retain model)?
  • 2
    Continual Unlearning: Is the method robust to sequential deletion requests without incurring catastrophic forgetting?
  • 3
    Detectability & Capacity: Consistent with our theory, does the model defeat perplexity-based detectors while retaining general capabilities?
We compare against state-of-the-art full-parameter unlearning baselines. Table 3 below summarizes these baselines alongside other unlearning approaches.
Method Unlearn objective Retain objective Unlearn nature
GA [yao2024llmunlearning] ascent on unlearn set full information
GD [liu2022continual] ascent on unlearn set descent on retain set full information
KL-GA [hong2024dissect] ascent on unlearn set minimize $\mathrm{KL}\!\big(p_{\theta}^{\,r}\,\|\,p_{\theta_{0}}^{\,r}\big)$ full information
DPO [rafailov2023dpo] preference loss minimize $\mathrm{CE}$ or $\mathrm{KL}$ full information
Forgetting–MarI minimize $I(X_{\mathrm{MarI}};Z)$ minimize $\mathrm{KL}\!\big(p_{\theta}^{\,r}\,\|\,p_{\theta_{0}}^{\,r}\big)$ marginal information
Table 3: Comparison of LLM unlearning objectives. CE = cross-entropy; KL = Kullback–Leibler divergence. “Direct/Indirect” indicates whether the method explicitly penalizes marginal information (ours) or approximates it by balancing forget/retain signals.
4.1 Models, Datasets, and Splits
Protocol. Following maini2024tofutaskfictitiousunlearning, eldan2023approximate, we adopt a three-step evaluation: (i) fine-tune on $\mathcal{D}_{u}\!\cup\!\mathcal{D}_{r}$ to obtain a full finetune baseline; (ii) apply each unlearning method to remove the influence of $\mathcal{D}_{u}$ to obtain unlearn results, denoted by the name of corresponding unlearn methods; (iii) restart the fine-tune but only on $\mathcal{D}_{r}$ (never exposed to $\mathcal{D}_{u}$) to obtain an unlearn baseline. An ideal unlearning method should preserve retain/validation accuracy level similar to the unlearn baseline while lowering the unlearn accuracy to a similar level to the unlearn baseline. We report next-token accuracy on $\mathcal{D}_{r}$, $\mathcal{D}_{u}$, and a held-out validation set, and assess general capability with the Eleuther’s LM Evaluation Harness [eval-harness].
Datasets. We evaluate Forgetting–MarI on two mid-scale LMs, GPT-2 Large (774M) and Llama-3.2-1B, and two text domains with distinct genres and pretraining prevalence: (i) Harry Potter and the Prisoner of Azkaban [rowling2013harry], likely present in pretraining; and (ii) Careless People: A Cautionary Tale of Power, Greed, and Lost Idealism [wynn2025careless], published after Llama’s release and thus unlikely to appear in its pretraining.
Splits. For Harry Potter (HP), we designate 10% of sentences as $\mathcal{D}_{u}$ and the remaining 90% as $\mathcal{D}_{r}$ (cf. eldan2023approximate); validation uses excerpts from another book in the series (Harry Potter and the Sorcerer’s Stone [rowling1999harry]). For Careless People (CP), we consider: (i) a correlated split with contiguous 50/50 spans for $\mathcal{D}_{u}/\mathcal{D}_{r}$; and (ii) an uncorrelated split where $\mathcal{D}_{u}$ comprises 2025 Reddit stories (post-release) and $\mathcal{D}_{r}$ is 50% of the book. Reddit stories are used as validation in the first setting and the remaining 50% of the book is used as validation in the second setting.
Parameter tuning & ablation. For each method, we stop training once validation accuracy drops by more than 3% from its initial value (to prevent general-utility degradation). Hyperparameter sweeps, ablations, and full training trajectories are provided in Appendix D.4.
Implementation details for hardware specifications and runtime information can be found in Appendix D.1.
4.2 Accuracy Trade-off & general capability
Figure 3: Left panels summarize the next-token accuracies on retain/unlearn/validation whereas the right panels summarize the general-capability on various benchmarks. Top row shows the results from Llama-3.2-1B on Careless People (correlated split), where as bottom row shows the results from GPT-2 Large on Harry Potter. Each method is reported at its best $\lambda$ and training epoch. On the left panels, an ideal method should match the unlearn baseline on retain/unlearn/validation accuracy on the left panels. On the right, better methods should achieve higher accuracy on ARC-E, HellaSwag, PIQA, MMLU and lower on WikiText perplexity test. Star indicates the best performer on that test.
Below, unlearn baseline is the model fine-tuned on retain set only, which serves as a retrain from scratch model.
Token-level utility (left panels). Across HP and CP, an ideal unlearning method should achieve similar retain/unlearn/validation accuracy as the unlearn baseline. Forgetting–MarI closely matches the unlearn baseline by achieving similar retain/unlearn/validation accuracy in both HP and CP experiments. In comparison, we observe unstable behaviors from methods with full-information unlearning nature: (i) Careless People: GD and DPO both overtrain on $\mathcal{D}_{r}$, while KL-GA struggles to remove the information in $\mathcal{D}_{u}$ while maintaining performance on $\mathcal{D}_{r}$. Moreover, all other methods show a gradual degradation on the validation accuracy. (The information of uncorrelated retain/unlearn split experiment is in Appendix D.) (ii) HP: GD and DPO are less overtrain on $\mathcal{D}_{r}$ but either over- or under-unlearn compared to the unlearn baseline. KL-GA still struggles to remove the information in $\mathcal{D}_{u}$ while maintaining performance on $\mathcal{D}_{r}$.
General capability (right panels). We compare the unlearned models against the finetuned baseline, the unlearn baseline, and the original GPT-2 (before finetuning) on ARC-E, PIQA, HellaSwag, WikiText, and MMLU [eval-harness].
  • Llama on Careless People. Forgetting–MarI attains the best results on HellaSwag and MMLU and ranks second on PIQA and ARC-E (DPO is first there but fails to unlearn effectively). It is also the only method that largely matches the full finetune baseline, which implies no model capacity degradation, because unlearning starts from full finetune baseline. Furthermore, notice that full finetuned baseline has lower general capacity than the unlearn baseline. That suggests a better finetuning could help Forgetting-MarI match the unlearn baseline across all benchmarks.
  • GPT-2 on Harry Potter. Forgetting–MarI is best on PIQA and WikiText, second on ARC-E, and slightly behind KL-GA/GD on HellaSwag. Furthermore, we note that high scores on general benchmarks do not guarantee meaningful utility: For instance, GA can outperform all other methods on MMLU despite lacking any utility preservation (and exhibiting very high WikiText perplexity which is consistent with theory). Thus, evaluating perplexity or token-level confidence is essential to avoid confounding factors.
In summary, Forgetting–MarI delivers targeted forgetting while largely preserving general capability, outperforming full-information baselines on the retain/unlearn/validation trade-off. See complete benchmark tables for experiment results on ARC-E, PIQA, MMLU, HellaSwag, and WikiText in Appendix D.5 (Tables 5 and 6).
4.3 Continual unlearning with sequential deletion requests
Setup.
For continual unlearning, we adopt a multi-stage protocol: we partition the total forget set into three disjoint subsets $\mathcal{D}_{u}=\mathcal{D}_{u,1}\cup\mathcal{D}_{u,2}\cup\mathcal{D}_{u,3}$. Starting from a baseline model fine-tuned on the full union $(\mathcal{D}_{r}\cup\mathcal{D}_{u})$, we perform three sequential unlearning steps: at step $t$, we unlearn $\mathcal{D}_{u,t}$ from the model resulting from step t1t-1. The subsets are defined as follows: (1) Harry Potter (Character-based): To emulate user-level unlearning, we build an HP “forget-characters” benchmark: we assign sentences to characters via alias matching (e.g., “Hermione”, “Hermione Granger”, “Granger” $\rightarrow$ Hermione). We target three characters sequentially: $\mathcal{D}_{u,1}$ (Hermione) $\to$ $\mathcal{D}_{u,2}$ (Snape) $\to$ $\mathcal{D}_{u,3}$ (Ron). The retain set $\mathcal{D}_{r}$ includes all “other” sentences, all retain-character sentences (e.g., Harry, Dumbledore), and sentences of not-yet-unlearned characters. (2) Careless People (Random-split): We randomly partition the designated forget set into three equal, disjoint subsets to form $\mathcal{D}_{u,1}$, $\mathcal{D}_{u,2}$, and $\mathcal{D}_{u,3}$. The retain set $\mathcal{D}_{r}$ consists of the remaining text from the book.
(a) GPT2-LG on Harry Potter (HP).
(b) Llama-3.2-1B on Careless People (CP).
Figure 4: Continual unlearning across models and datasets. Rows show methods and sequential steps; columns report retain/validation accuracies, performance on chunk-specific content, and general capability. Darker color indicates higher accuracy. For general capabilities (ARC-E, PIQA and WikiText): higher accuracies/lower perplexities and stable performance across steps indicates successful knowledge preservation. For unlearned content: ideal pattern shows accuracy drop after removal, with persistent forgetting in subsequent steps.
Results. Figures 4 summarize GPT-2/HP and Llama/CP results, respectively:
  • GPT-2 on HP. Forgetting–MarI is the only method that remains robust across retain and validation accuracy, preserves forgetting on previously unlearned sets, and sustains general capability. By contrast, KL-GA tends to relearn previously forgotten content and substantially degrades WikiText performance; DPO fails to effectively unlearn; GD overshoots on $\mathcal{D}_{r}$ and also loses general capability (WikiText).
  • Llama on CP. Forgetting–MarI again exhibits the most consistent behavior: retain/validation accuracy and previously unlearned sets remain stable, and ARC-E/PIQA show minimal drift. In comparison, KL-GA quickly over-unlearns, lowering retain/validation and prior-unlearn accuracies; DPO again fails to unlearn effectively; GD overshoots on $\mathcal{D}_{r}$, reduces validation accuracy, and shows deteriorating ARC-E/PIQA performance.
Conclusion. Forgetting–MarI delivers robust sequential unlearning: it preserves utility, maintains prior forgetting, and sustains general capability across steps, outperforming full-information baselines in both performance and stability.
4.4 Detector evaluation: Empirical verification of theoretical guarantees